# Fibonacci-Modulation-Induced Multiple Topological Anderson Insulators

### v1

@@ Acknowledgements.
Supplementary Materials for ”Fibonacci-Modulation-Induced Multiple Topological Anderson Insulators”
This Supplementary Material provides detailed explanations of the scattering-matrix method used to calculate the $\mathcal{Z}_{2}$ topological quantum number $Q$, the derivation of topological phase boundaries, an analysis of an off-diagonal Fibonacci-modulated SOC chain, and the experimental realization.
I. Scattering-matrix method
To determine the topological properties of the quasiperiodic-modulated SOC chain, we utilize the scattering matrix $S$, which relates the incoming and outgoing wave amplitudes at the Fermi level [1]:
$S=\begin{pmatrix}\tilde{R}_{\leftarrow}&\tilde{T}_{\leftarrow}\\ \tilde{T}_{\rightarrow}&\tilde{R}_{\rightarrow}\\ \end{pmatrix},$ (S1)
$Q=\mathrm{sgn}\left[\mathrm{Det}\left(\tilde{R}_{\leftarrow}\right)\right]=\mathrm{sgn}\left[\mathrm{Det}\left(\tilde{R}_{\rightarrow}\right)\right],$ (S2)
where $\mathrm{sgn}[\cdots]$ denotes the sign function. A value of $Q=1$ indicates a topologically trivial phase, while $Q=-1$ signifies a topologically nontrivial phase. To account for the effects of disorder, we define the disorder-averaged topological invariant $\overline{Q}$, calculated over $N_{c}$ disorder realizations. In our analysis, $N_{c}$ configurations are sampled for different values of $\phi\in[0,2\pi]$.
$\begin{pmatrix}\mathcal{R}_{i+1,i}^{\dagger}\Psi_{i}\\ \Psi_{i+1}\end{pmatrix}=\tilde{B}_{i}\begin{pmatrix}\mathcal{R}_{i,i-1}^{\dagger}\Psi_{i-1}\\ \Psi_{i}\end{pmatrix}$ (S3)
where
$\tilde{B}_{i}=\begin{pmatrix}0&\mathcal{R}_{i+1,i}^{\dagger}\\ -\mathcal{R}_{i,i+1}^{-1}&-\mathcal{R}_{i,i+1}^{-1}\mathcal{U}_{i}\\ \end{pmatrix}.$ (S4)
$B_{i}=U^{\dagger}\tilde{B}_{i}U,$ (S5)
where
$U=\frac{1}{\sqrt{2}}\begin{pmatrix}I&I\\ iI&-iI\end{pmatrix},$ (S6)
and $I$ is a $2\times 2$ identity matrix. In this base, the reflection $(\tilde{R}_{\leftarrow},\tilde{R}_{\rightarrow})$ and transmission $(\tilde{T}_{\leftarrow},\tilde{T}_{\rightarrow})$ matrices can be determined via the relations:
$\begin{pmatrix}\tilde{T}_{\rightarrow}\\ 0\end{pmatrix}=B\begin{pmatrix}I\\ \tilde{R}_{\leftarrow}\end{pmatrix},\quad\begin{pmatrix}\tilde{R}_{\rightarrow}\\ I\end{pmatrix}=B\begin{pmatrix}0\\ \tilde{T}_{\leftarrow}\end{pmatrix},$ (S7)
where $B=B_{L}B_{L-1}\ldots B_{1}$.
II. Derivation of topological phase boundaries
@@
In the topologically nontrivial regime, the zero edge modes emerge with a finite localization length. Conversely, in the trivial regime, these edge modes disappear, giving way to bulk states characterized by a divergence in localization length. To analytically determine the topological phase boundaries, we examine the localization length of the zero modes. The zero mode Schrödinger equation for the Fibonacci-modulated model, $H|\Psi\rangle=0$, is expressed as:
@@
where $\psi_{i\beta}$ represents the probability amplitude of the zero mode for spin $\beta=\{\uparrow,\downarrow\}$ at the $i$th lattice, and $U_{i}=[\lambda\mathrm{sgn}(\cos{(2\pi\alpha i+\phi)}-\cos{(\pi\alpha)})+M]$. Applying a local transformation $\phi^{+}_{i}=\psi_{i\uparrow}+\psi_{i\downarrow}$ and $\phi^{-}_{i}=\psi_{i\uparrow}-\psi_{i\downarrow}$ [2], the equation decouple into two independent equations of similar form:
@@
Focusing on one decoupled equation, the evolution can be written in terms of the transfer matrix:
@@
where the transfer matrix $T_{i}$ is:
@@
$|\mathrm{Tr}(T)|-\mathrm{Det}(T)=1.$ (S15)
III. Off-diagonal-Fibonacci-modulated SOC chain
Figure S1: (a)Topological phase diagram of an off-diagonal-Fibonacci-modulated SOC chain in $\Delta-M$ plane with $L=610$, $\alpha=F_{13}/F_{14}$, and $t_{so}=0.25$. The blue dashed lines represent the topological phase boundaries numerically obtained by Eq. (S15). (b) Two disorder-averaged energies $\overline{E}_{L}$ and $\overline{E}_{L+1}$ at the center of the spectrum and the disorder-averaged topological number $\overline{Q}$ as a function of $\Delta$ under OBC with $t_{so}=0.25$ and $M=-4$. (c) The number of times that the TAI phase emerges as a function of $t_{so}$ with $M=-4$. Here, all the data are averaged by $N_{c}=50$ disorder realizations.
Figure S2: (a)Self-similarities of energy spectrum as a function of $\Delta$ with $t_{so}=0$. Two central energy levels $E_{L}$ and $E_{L+1}$ as a function of $\Delta$ with (b) $t_{so}=0.2$, (c) $0.3$, and (d) $0.9$. Here, $M=-4$, $\phi=0$, and $L=610$.
Figure S3: Fractal dimensions $D_{n}$ for different system sizes under PBC for (a) $\Delta=2.4$ and (b) $\Delta=3$, respectively. (c) The scaling of the mean fractal dimensions $\langle D\rangle$ for various $\Delta$ values. Here, $t_{so}=0.25$, $M=-4$, and $\phi=0$.
@@ III. Off-diagonal-Fibonacci-modulated SOC chain
where the $\pm$ corresponds to the hopping along the ±$\pm\hat{x}$ direction. The diagonal terms of $\mathcal{R}$ describe the spin-conserved hopping with the amplitude $t_{0}$, modulated by the Fibonacci sequence with strength $\Delta$. The off-diagonal terms represent spin-flip hopping between nearest neighbors with an amplitude $\mathcal{R}$. We set $t_{0}=1$ as the unit of energy throughout this discussion.
We calculated the topological phase diagram in the $\Delta-M$ plane, as shown in Fig. S1(a), using $L=610$, $t_{so}=0.25$, and $N_{c}=50$ disorder realizations. The dashed lines represent the topological phase boundaries, numerically determined by the condition $|\mathrm{Tr}(T)|-\mathrm{Det}(T)=1$, where $T=\prod_{i=1}^{L}T_{i}$ and
@@
where, $\xi_{i}^{(\pm)}=t_{so}\pm\left[t_{0}-\frac{\Delta}{2}\mathrm{sgn}\left[\cos(2\pi\alpha i+\phi)-\cos(\pi\alpha)\right]+\frac{\Delta}{2}\right]$. The results show that the topologically nontrivial phase is robust against weak modulation for $20<\Delta<$ 2 but transitions to a trivial phase as the modulation strength increases. Notably, when $M=0$, the system remains in the topological phase regardless of $\Delta$. For $3.052<|M|<$ 3.05, the system undergoes a topological phase transition from trivial to nontrivial and back to trivial with increasing $\Delta$, demonstrating the emergence of the TAI phase. When $|M|\geq 3.05$, the system exhibits multiple transitions into the TAI phase before ultimately reverting to a trivial phase at higher $\Delta$. Figure S1(b) shows the disorder-averaged energies of the two center modes, $\overline{E}_{L}$ and $\overline{E}_{L+1}$, alongside the $\mathcal{Z}_{2}$ topological number $\overline{Q}$ as a function of the modulation strength $\Delta$, with $M=-4$ and $t_{so}=0.25$. The results indicate that the moderate modulation $\Delta$ induces changes in the nontrivial invariant within the regions $\Delta\in[2.28,2.49]\cup[2.66,2.86]\cup[3.22,3.44]$, where disorder-averaged zero modes are also observed. For $\Delta>3.44$, the TAI phase vanishes. In Fig. S1(c), we show the number of times the TAI phase emerges, $\overline{g}$, as a function of $t_{so}$ for $M=-4$. Similar to the on-site modulated case, $\overline{g}$ in the off-diagonal-modulated case increases as $t_{so}$ decreases. To understand this phenomenon, we numerically calculate the energy spectrum of two decoupled off-diagonal-Fibonacci-modulated chains without the SOC term ($t_{so}=0$) as a function of $\Delta$, shown in Fig. S2(a) for $M=-4$ and $\phi=0$. Within the range $\Delta\in(1.16,2.41)$, the middle bands overlap, forming three distinct clusters of band-crossing regions separated by two band gaps near zero energy. These clusters further split into sub-clusters, creasing a self-similar fractal structure in the thermodynamic limit, as seen in Fig. S2(a). As the SOC strength $t_{so}$ increases, these band gaps, with magnitudes comparable to $t_{so}$, progressively close and reopen, leading to the emergence of nontrivial zero-energy modes in these gaps, as depicted in Fig. S2(b). This suggests an infinite number of TAI phases emerge in the small $t_{so}$ limit and in the thermodynamic limit, forming a topologically fractal structure. As $t_{so}$ increases further, adjacent topologically nontrivial regions merge, reducing the number of distinct TAI phases, as illustrated in Figs. S2(b) and S2(c). These findings confirm that an off-diagonal Fibonacci-modulated SOC chain supports multiple TAI phases, with their occurrence strongly dependent on $t_{so}$.
@@
Figures S3(a) and S3(b) show the fractal dimensions for different system sizes under periodic boundary condition (PBC) with $\phi=0$ for $\Delta=2.4$ and $\Delta=3$, respectively. For both topologically trivial and nontrivial phases, the fractal dimensions of nearly all states are system-size independent, deviating from the extremes of 00 and $1$. The scaling behavior of the mean fractal dimension $\langle D\rangle$ for various $\Delta$ values, in both topologically trivial ($\Delta=2.4$, $2.8$, and $3.3$) and nontrivial ($\Delta=2$, $2.6$, $3$, and $3.6$) phases, is shown in Fig. S3(c). In the thermodynamic limit, $\langle D\rangle$ approaches values that deviate from both 00 and $1$ in both phases. Our results suggest that the emergence of the TAI phase does not affect the localization properties of the systme, and the states remain multifractal.
IV. Experimental realization
Figure S4: Illustration of experimental scheme. (a) A quasi-1D BEC is illuminated by a pair of counterpropagating laser beams: one with a fixed frequency $\omega$ and the other containing multifrequency components $\omega_{j}$. (b) The lasers, far detuned from the atomic transition, drive a series of engineered two-photon Bragg transitions. These transitions couple different momentum states within the same ground-state hyperfine manifold (blue and yellow arrows) or across different ground-state hyperfine manifolds (black arrows), with an increment of $2\hbar\tilde{k}$, where $\tilde{k}=2\pi/\tilde{\lambda}$ with $\tilde{\lambda}$ being the wave length of the lasers. The solid and dashed arrows represent the processes of photon absorption and emission, respectively.
Our Fibonacci-modulated SOC model can be implemented along the momentum lattice in a cold atom system using a Bose-Einstein condensate (BEC), as illustrated in Fig. S4. In this setup, the spin and lattice-site degrees of freedom are encoded in the two ground-state hyperfine manifolds of selected atoms, such as 87Rb [3]. Specifically, the hyperfine state $F=1$ represents spin-down, while $F=2$ represents spin-up. A pair of counter-propagating laser beams is employed: the incoming laser beam has a fixed frequency, while the acousto-optic modulator in the reflected laser beam generate multi-frequency components. These laser beams drive a series of two-photon Bragg transitions between adjacent discrete momentum states. Transitions within the same hyperfine manifold enable nearest-neighbor couplings with the same spin, corresponding to the diagonal terms of $\mathrm{R}$. In contrast, transitions between different hyperfine manifolds facilitate nearest-neighbor couplings with opposite spins, corresponding to the off-diagonal terms of $\mathrm{R}$. By precisely adjusting the amplitude of each multi-frequency component, all couplings can be individually controlled. Additionally, the on-site potentials $\mathcal{U}_{i}$ can be tuned via the detuning of the two-photon Bragg transitions.

### v3

@@ Acknowledgements
Author contributions statement
Z.H.X. and R.J.J. conceived and designed the project. R.J.J. performed the numerical simulations. Z.H.X. provided the explanation of the numerical results. All authors contributed to the discussion of the results and wrote the paper.
Competing Interests
The authors declare no competing interests.
@@
Supplementary Information for ”Fibonacci-Modulation-Induced Multiple Topological Anderson Insulators”
This Supplementary Information details the procedures for generating generalized $\kappa$-Fibonacci-modulated chain, presents the derivation of topological phase boundaries, analyzes off-diagonal Fibonacci-modulated SOC chains, and discusses the impact of modulation frequency on topological properties.
Supplementary Note 1: Generate a generalized $\kappa$-Fibonacci-modulated chain
Three methods are commonly employed to generate a $\kappa$-Fibonacci-modulated chain: the substitution method, the cut-and-project method, and the characteristic function method. In the substitution method, the generalized $\kappa$-Fibonacci substitution rule $\tilde{\sigma}$ acts on the two letters $A$ and $B$ as follows:
$\tilde{\sigma}:\begin{cases}A\to A^{\kappa}B,\\ B\to A,\end{cases}$ (S1)
where $\kappa=1,2,3,\cdots$ corresponds to different metallic mean sequences. By repeatedly applying this substitution on the initial letter $B$, one generates a sequence of words $C_{l}=\tilde{\sigma}^{l}(B)$ of increasing length. These finite sequences serve as approximates to the infinite generalized $\kappa$-Fibonacci chain, which is recovered in the limit $l\to\infty$. The inflation matrix associated with this substitution has eigenvalues that satisfy the characteristic equation $\lambda^{\prime 2}-\kappa\lambda^{\prime}-1=0$. The Perron-Frobenius eigenvalues for several cases are: $\lambda^{\prime}_{g}=(\sqrt{5}+1)/2$ for $\kappa=1$ (golden mean), $\lambda^{\prime}_{s}=\sqrt{2}+1$ for $\kappa=2$ (silver mean), and $\lambda^{\prime}_{b}=(\sqrt{13}+3)/2$ for $\kappa=3$ (bronze mean). The modulation frequency $\alpha$ of the Fibonacci-modulated chain is chosen as the inverse of the Perron-Frobenius eigenvalue, i.e., $\alpha=1/\lambda^{\prime}$. For numerical calculations, $\alpha$ is approximated by the rational number $\alpha=F_{\nu-1}/F_{\nu}$, where the generalized Fibonacci numbers $F_{\nu}$ are defined recursively by $F_{\nu+1}=F_{\nu-1}+\kappa F_{v}$ with initial conditions $F_{0}=F_{1}=1$.
The generalized $\kappa$-Fibonacci chain can also be constructed using the cut-and-project method, which projects a 2D periodic lattice onto a 1D quasicrystal. As illustrated in Fig. S1, the parent system is a 2D square lattice, and the quasicrystalline chain is obtained by projecting lattice points onto the physical axis $\tilde{x}$, which is oriented at an angle $\tilde{\theta}$ such that $\tan{\tilde{\theta}}=\alpha$. A periodic lattice point is selected if it lies within a green strip of width equal to one unit cell of the 2D lattice. Upon projection onto the $\tilde{x}$ axis, the horizontal and vertical bonds of the square lattice project onto the red and blue intervals, respectively. The resulting spacings between adjacent points on the projected chain—represented by these red and blue intervals—can take two values, $\cos{\tilde{\theta}}=1/\sqrt{1+\alpha^{2}}$ and $\sin{\tilde{\theta}}=\alpha/\sqrt{1+\alpha^{2}}$. These two lengths corresponds to the $A$ and $B$ tiles in the generalized $\kappa$-Fibonacci sequence, respectively.
Figure S1: Schema of the cut-and project method. Selected points of a 2D square lattice are projected onto the physical axis $\tilde{x}$, giving the quasiperiodic sequence of A(red) and B(blue) tiles.
The third approach for generating the generalized $\kappa$-Fibonacci quasicrystal is the characteristic function method. Kraus and Zilberberg Zilberberg2012fulu proposed a scheme that bridges the Fibonacci and AA models. In this method, the $i$-th letter of the $\kappa$-Fibonacci chain is determined by the characteristic function $\chi_{i}$, defined by
$\chi_{i}=\mathrm{sgn}[\cos(2\pi\alpha i+\phi)-\cos(\pi\alpha)],$ (S2)
where $\phi$ is an arbitrary phase. According to this definition, $\chi_{i}=+1$ corresponds to the letter $A$, while $\chi_{i}=-1$ corresponds to the letter $B$.
Supplementary Note 2: Derivation of topological phase boundaries
@@
In the topologically nontrivial regime, the zero-energy edge modes emerge with a finite localization length. Conversely, in the trivial regime, these edge modes vanish, giving way to bulk states characterized by a divergence in localization length. To analytically determine the topological phase boundaries, we examine the localization length of the zero modes. The zero mode Schrödinger equation for the Fibonacci-modulated model, $H|\Psi\rangle=0$, is expressed as:
@@
where $\psi_{i\beta}$ represents the amplitude of the zero mode with spin $\beta=\{\uparrow,\downarrow\}$ at site $i$, and $U_{i}=[\lambda\mathrm{sgn}(\cos{(2\pi\alpha i+\phi)}-\cos{(\pi\alpha)})+M]$. By introducing the local basis transformation $\phi^{+}_{i}=\psi_{i\uparrow}+\psi_{i\downarrow}$ and $\phi^{-}_{i}=\psi_{i\uparrow}-\psi_{i\downarrow}$, the coupled equations decouple into two independent equations of similar form:
@@
Focusing on one of these decoupled equations, the evolution can be rewritten in terms of a transfer matrix:
@@
where the transfer matrix $T_{i}$ is
@@
where $T\equiv\prod_{i=1}^{L}T_{i}$ is the total transfer matrix. Here, $\|\cdot\|$ denotes the matrix Euclidean norm, which is defined as the largest singular value of the matrix. For a normal matrix, such as the transfer matrix considered here, the Euclidean norm is equivalent to the absolute value of its largest eigenvalue. For the $2\times 2$ matrix $T$, the eigenvalues are
$\lambda_{1}=\frac{\mathrm{Tr}(T)}{2}+\sqrt{\left[\frac{\mathrm{Tr}(T)}{2}\right]^{2}-\mathrm{Det}(T)},\quad\lambda_{2}=\frac{\mathrm{Tr}(T)}{2}-\sqrt{\left[\frac{\mathrm{Tr}(T)}{2}\right]^{2}-\mathrm{Det}(T)}.$ (S10)
The phase boundary between topological and trivial regimes is defined by the condition $\tilde{\lambda}^{-1}=0$, which requires that the largest eigenvalue of $T$ satisfies $\left|\max\{\lambda_{1},\lambda_{2}\}\right|=1$. Solving this condition yields the phase boundary equation: $|\mathrm{Tr}(T)|-\mathrm{Det}(T)=1$, which is shown in the main text.
Supplementary Note 3: Off-diagonal-Fibonacci-modulated SOC chain
@@
where the $\pm$ corresponds to the hopping along the $\pm\hat{x}$ direction. The diagonal terms of $\mathcal{R}$ describe the spin-conserved hopping with the amplitude $t_{0}$, modulated by the Fibonacci sequence with strength $\Delta$. The off-diagonal terms represent spin-flip hopping between nearest neighbors with an amplitude $\mathcal{R}$. We set $t_{0}=1$ as the unit of energy throughout this discussion. We calculated the topological phase diagram in the $\Delta-M$ plane, as shown in Fig. S2(a), using $L=610$, $t_{so}=0.25$, and $N_{c}=50$ disorder realizations. The dashed lines represent the topological phase boundaries, numerically determined by the condition $|\mathrm{Tr}(T)|-\mathrm{Det}(T)=1$, where $T=\prod_{i=1}^{L}T_{i}$ and
@@
Figure S2: (a)Topological phase diagram of an off-diagonal-Fibonacci-modulated SOC chain in $\Delta-M$ plane with $L=610$, $\alpha=F_{13}/F_{14}$, and $t_{so}=0.25$. The blue dashed lines represent the topological phase boundaries numerically obtained by Eq. (3) in the main text. (b) Two disorder-averaged energies $\overline{E}_{L}$ and $\overline{E}_{L+1}$ at the center of the spectrum and the disorder-averaged topological number $\overline{Q}$ as a function of $\Delta$ under OBC with $t_{so}=0.25$ and M=4M=-4. (c) The number of times that the TAI phase emerges as a function of $t_{so}$ with M=4M=-4. Here, all the data are averaged by $N_{c}=50$ disorder realizations.
Figure S3: (a)Self-similarities of energy spectrum as a function of $\Delta$ with $t_{so}=0$. Two central energy levels $E_{L}$ and $E_{L+1}$ as a function of $\Delta$ with (b) $t_{so}=0.2$, (c) $0.3$, and (d) $0.9$. Here, M=4M=-4, $\phi=0$, and $L=610$.
Figure S4: Fractal dimensions $D_{n}$ for different system sizes under PBC for (a) $\Delta=2.4$ and (b) $\Delta=3$, respectively. (c) The scaling of the mean fractal dimensions $\langle D\rangle$ for various $\Delta$ values. Here, $t_{so}=0.25$, M=4M=-4, and $\phi=0$.
@@
where, $\xi_{i}^{(\pm)}=t_{so}\pm\left[t_{0}-\frac{\Delta}{2}\mathrm{sgn}\left[\cos(2\pi\alpha i+\phi)-\cos(\pi\alpha)\right]+\frac{\Delta}{2}\right]$. The results show that the topologically nontrivial phase is robust against weak modulation for $0<\Delta<2$ but transitions to a trivial phase as the modulation strength increases. Notably, when $M=0$, the system remains in the topological phase regardless of $\Delta$. For $2<|M|<3.05$, the system undergoes a topological phase transition from trivial to nontrivial and back to trivial with increasing $\Delta$, demonstrating the emergence of the TAI phase. When $|M|\geq 3.05$, the system exhibits multiple transitions into the TAI phase before ultimately reverting to a trivial phase at higher $\Delta$. Figure S2(b) shows the disorder-averaged energies of the two center modes, $\overline{E}_{L}$ and $\overline{E}_{L+1}$, alongside the $\mathcal{Z}_{2}$ topological number $\overline{Q}$ as a function of the modulation strength $\Delta$, with M=4M=-4 and $t_{so}=0.25$. The results indicate that the moderate modulation $\Delta$ induces changes in the nontrivial invariant within the regions $\Delta\in[2.28,2.49]\cup[2.66,2.86]\cup[3.22,3.44]$, where disorder-averaged zero modes are also observed. For $\Delta>3.44$, the TAI phase vanishes. In Fig. S2(c), we show the number of times the TAI phase emerges, $\overline{g}$, as a function of $t_{so}$ for M=4M=-4. Similar to the on-site modulated case, $\overline{g}$ in the off-diagonal-modulated case increases as $t_{so}$ decreases. To understand this phenomenon, we numerically calculate the energy spectrum of two decoupled off-diagonal-Fibonacci-modulated chains without the SOC term ($t_{so}=0$) as a function of $\Delta$, shown in Fig. S3(a) for M=4M=-4 and $\phi=0$. Within the range $\Delta\in(1.16,2.41)$, the middle bands overlap, forming three distinct clusters of band-crossing regions separated by two band gaps near zero energy. These clusters further split into sub-clusters, creasing a self-similar fractal structure in the thermodynamic limit, as seen in Fig. S3(a). As the SOC strength $t_{so}$ increases, these band gaps, with magnitudes comparable to $t_{so}$, progressively close and reopen, leading to the emergence of nontrivial zero-energy modes in these gaps, as depicted in Fig. S3(b). This suggests an infinite number of TAI phases emerge in the small $t_{so}$ limit and in the thermodynamic limit, forming a topologically fractal structure. As $t_{so}$ increases further, adjacent topologically nontrivial regions merge, reducing the number of distinct TAI phases, as illustrated in Figs. S3(c) and S3(d). These findings confirm that an off-diagonal Fibonacci-modulated SOC chain supports multiple TAI phases, with their occurrence strongly dependent on $t_{so}$.
@@
Figures S4(a) and S4(b) show the fractal dimensions for different system sizes under periodic boundary condition (PBC) with $\phi=0$ for $\Delta=2.4$ and $\Delta=3$, respectively. For both topologically trivial and nontrivial phases, the fractal dimensions of nearly all states are system-size independent, deviating from the extremes of 0 and $1$. The scaling behavior of the mean fractal dimension $\langle D_{r}\rangle$ for various $\Delta$ values, in both topologically trivial ($\Delta=2.4$, $2.8$, and $3.3$) and nontrivial ($\Delta=2$, $2.6$, 33, and $3.6$) phases, is shown in Fig. S4(c). In the thermodynamic limit, $\langle D_{r}\rangle$ approaches values that deviate from both 0 and $1$ in both phases. Our results suggest that the emergence of the TAI phase does not affect the localization properties of the system, and the states remain multifractal.
Supplementary Note 4: Influence of modulation frequency on the emergence of multiple reentrant TAI phases
Figure S5: (a) The disorder-averaged topological phase diagram of the system in $M-\lambda$ plane with $t_{so}=0.5$. The blue dashed lines correspond to the topological phase boundaries determined by Eq.(3) in the main text. (b) The number of times the TAI phases emerge, denoted as $\overline{g}$, as a function of $t_{so}$ with $M=4$. Here, all the data are averaged by $N_{c}=50$ disorder realizations. (c) Self-similarities of energy spectrum as a function of $\lambda$ with $t_{so}=0$. Two central energy levels $E_{L}$ and $E_{L+1}$ as a function of $\lambda$ with (d) $t_{so}=0.25$, (e) $0.5$, and (f) $0.9$. Here, $\kappa=2$, $M=4$, $\phi=0$, and $L=577$.
Figure S6: (a) The disorder-averaged topological phase diagram of the system in $M-\lambda$ plane with $t_{so}=0.5$. The blue dashed lines correspond to the topological phase boundaries determined by Eq.(3) in the main text. (b) The number of times the TAI phases emerge, denoted as $\overline{g}$, as a function of $t_{so}$ with $M=4$. Here, all the data are averaged by $N_{c}=50$ disorder realizations. (c) Self-similarities of energy spectrum as a function of $\lambda$ with $t_{so}=0$. Two central energy levels $E_{L}$ and $E_{L+1}$ as a function of $\lambda$ with (d) $t_{so}=0.27$, (e) $0.5$, and (f) $0.9$. Here, $\kappa=3$, $M=4$, $\phi=0$, and $L=469$.
To investigate the influence of modulation frequency on the emergence of multiple TAIs, we extend our analysis to other metallic mean numbers ($\kappa=2$ and $3$), in contrast to the main text where $\alpha$ is chosen as the golden mean $(\sqrt{5}-1)/2$, corresponding to $\kappa=1$.
For the silver mean ($\kappa=2$), Fig. S5(a) displays the disorder-averaged topological phase diagram for $t_{so}=0.5$ as a function of $M$ and $\lambda$. This diagram is obtained by numerically averaging $\mathcal{Z}_{2}$ invariant $\overline{Q}$ over $N_{c}$ disorder realizations for various values of $\phi\in[0,2\pi)$, with phase boundaries determined according to Eq.(3) in the main text. For $M>2$, the system exhibits multiple distinct transitions into the TAI phase, eventually reverting to a trivial phase as $\lambda$ increases further. We also examine how the number of TAI phases varies with $t_{so}$ at $M=4$, as shown in Fig. S5(b). It is evident that the average number of emergent TAI phases $\overline{g}$ increases as $t_{so}$ decreases. Following the approach in the main text, we further analyze the energy spectrum of two decoupled Fibonacci-modulated chains with $\kappa=2$ in the absence of the SOC term ($t_{so}=0$), as depicted in Fig. S5(c) for $\phi=0$. Here, a self-similar fractal structure is observed. Upon introducing a finite SOC strength $t_{so}$, band gaps comparable in size to $t_{so}$ begin to close and reopen, indicating the emergence of nontrivial zero-energy modes, as illustrated in Fig. S5(d). As $t_{so}$ increases further, adjacent topologically nontrivial regions merge, reducing the number of distinct TAI phases. This transition, shown in Figs. S5(e) and S5(f), marks the evolution from a fractal topological structure to a single, more conventional one.
For the bronze mean ($\kappa=3$), Fig. S6(a) shows the disorder-averaged topological phase diagram for $t_{so}=0.5$ as a function of $M$ and $\lambda$, obtained using the same methodology. The phase boundaries are again determined by Eq.(3) in the main text. For $M>3.1$, the system undergoes multiple distinct transitions into the TAI phase, eventually returning to a trivial phase as $\lambda$ increases further. Figure S6(b) presents the dependence of the number of emergent TAI phases on $t_{so}$ at $M=4$. We find that $\overline{g}$ increases as $t_{so}$ decreases, with a growth rate even faster than in the golden mean and silver mean cases. Using the same approach, we analyze the energy spectrum of two decoupled Fibonacci-modulated chains with $\kappa=3$ without the SOC term ($t_{so}=0$), as shown in Fig. S6(c) for $\phi=0$, and again observe a self-similar fractal structure. When the SOC strength $t_{so}$ is introduced, band gaps comparable in size to $t_{so}$ start to close and reopen, signaling the formation of nontrivial zero-energy modes, as depicted in Fig. S6(d). As $t_{so}$ increases further, neighboring topologically nontrivial regions merge, reducing the number of distinct TAI phases. This transition, illustrated in Figs. S6(e) and S6(f), signifies a transformation from a fractal topology to a more discrete structure.
Overall, our findings demonstrate that the emergence of multiple reentrant TAI phases is a universal and robust phenomenon across different types of Fibonacci modulation frequencies, consistent with the results presented in the main text. By employing various metallic mean modulations, such as the silver and bronze means, we consistently observe self-similar fractal structures in the energy spectra and multiple transitions between trivial and topological phases as the disorder strength varies. Notably, the modulation frequency not only determines the complexity of the fractal structure but also significantly influences the evolution of TAI phases. These results highlight the universal and robust nature of reentrant TAI behavior induced by fractal modulations, and underscore the critical role of modulation frequency in shaping the topological phase diagram of the system.